OpenStudy (anonymous):

What is the absolute value of the number?3 + i a:sqaureroot10 b:4 c:8 d:squareroot3

jimthompson5910 (jim_thompson5910):

Hint: \[\Large |a+bi| = \sqrt{a^2+b^2}\]

OpenStudy (anonymous):

4?!

jimthompson5910 (jim_thompson5910):

a = 3 and b = 1

jimthompson5910 (jim_thompson5910):

plug them in

jimthompson5910 (jim_thompson5910):

\[\Large |a+bi| = \sqrt{a^2+b^2}\] \[\Large |3+1i| = \sqrt{3^2+1^2}\] \[\Large |3+i| = ?\]

OpenStudy (anonymous):

ok i got 9 some how

jimthompson5910 (jim_thompson5910):

close

jimthompson5910 (jim_thompson5910):

3^2 is 9, but that's just part of it

OpenStudy (anonymous):

ok so now what

jimthompson5910 (jim_thompson5910):

simplify \[\Large \sqrt{3^2+1^2}\]

jimthompson5910 (jim_thompson5910):

what do you get

OpenStudy (anonymous):

sqaure root10?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

you are correct